# Hyper-Hermitian Weyl Double Copy

### v1

@@ Abstract
The self-dual double copy is further explored. In previous work, it has been shown that Hyper-Hermitian manifolds also have associated the self-dual gauge theories via Kerr-Schild double copy. The self-dual double copy is generalized in the structure of the kinematic algebra by changing the area-preserving diffeomorphisms algebra by the diffeomorphisms on a surface algebra. This gave rise to the hyper-Hermitian double copy in the Kerr-Schild approach. In the present article, we further study the hyper-Hermitian case using the Weyl double copy formalism. In particular, we have found solutions within this formalism for different hyper-Hermitian metrics. One of the main features is that there will be two Maxwell spinors and one of them is source-free and the other has a source current. This is compatible with the fact that in general the hyper-Hermitian spaces are not Ricci-flat.
@@ 1 Introduction
In the past, the interplay between different theoretical physics subjects has been a way of understanding physical phenomena and is a road to advancing in physics. For instance, AdS/CFT (or gauge/gravity) correspondence is a duality which (in its more traditional version) connects strong coupling supersymmetric quantum conformal gauge theory in dimension $d$ with classical gravity theory in $d+1$ dimensions (see for instance, [1]). Another relation between gauge and gravitational theories is the one of the double copy and the Bern-Carrasco-Johansson (BCJ) duality [2, 3, 4] which relates scattering amplitudes of non-abelian gauge theories in $d$ dimensions and scattering amplitudes of gravitational theories in the same dimension $d$ (for a complete review, see [5, 6]). The tree-level version has been proved and the loop extension is still at the level of conjecture.
@@
The idea of the double copy and the BCJ duality have been applied at the level of equations of motion of gauge and gravity theories. Thus it is promoted to the level of solutions of gauge and gravity theories. There is a correspondence between solutions of the equation of motion of both theories [7]. The original version of the classical double copy uses the Kerr-Schild ansatz which implies the existence of certain coordinate systems in which the metric of a gravity theory can be expressed as $g_{\mu\nu}=\eta_{\mu\nu}+\phi k_{\mu}k_{\nu}$. Many applications in diverse field theories were studied for this classical double copy approach, for instance, see refs. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Later another possibility to study the correspondence was through the use of Weyl spinors, which was named the Weyl double copy [23], Weyl double copy is reduced to Kerr-Schild double copy in type D spacetimes [24]. The Weyl double copy relates the curvature of gravity theory written in terms of Weyl spinors to the spinor field strength of a gauge field. After its proposal, Weyl double copy was discussed in different contexts in refs. [25, 26, 27, 28, 29]. This was the precursor to finding a more general description of the double copy in terms of twistor theory [30, 31, 32, 33, 34, 35]. Among the successes of the double copy is the one of considering it as an efficient tool for computing processes involving gravitational waves [27].
@@
In particular, the case for the self-dual sector of the Yang-Mills theories in four dimensions and its correspondence with the self-dual sector of Einstein’s theory was proposed and explored for the first time within the context of the double copy[36, 37, 38]. Integrability anomalies arising in the self-dual sectors of gravity and Yang-Mills theory are studied in [39]. This case is of particular relevance since both gauge and gravitational sectors can be described through a cubic Lagrangian giving rise to trivalent Feynman graphs. In the self-dual gauge case, one has the Leznov-Parks Lagrangian [40, 41] and in the gravitational self-dual sector, we have Plebański Lagrangian [42, 43]. In ref. [44] it was studied new cases for the heavenly double copies. There were considered integrable and non-integrable deformations of the self-dual double copy and generalizations of it as the hyper-Hermitian case. The double copy has been studied as well in the context of celestial amplitudes [45, 46, 47] and higher spin self-dual theories [48]. Some recent works regarding further developments of the self-dual double copy were discussed in some references [49, 50, 51, 52].
@@
In the present article, we continue the study started in [44] and we carry out a deeper study of the hyper-Hermitian double copy using the Weyl double copy. The hyper-Hermitian metrics have been studied previously in [53, 54, 55, 56, 57]. In [57], hyper-Hermitian complex metrics are determined through two holomorphic functions $\Theta_{A}(x^{a})$ (with $A=0,1$) of the coordinates on a complex four manifold. Moreover $\Theta_{A}(x^{a})$ satisfies a generalization of the Plebański second heavenly equation [58] that we term it the Dunajski heavenly equation. Indeed this generalized equation can be reduced to the second heavenly equation under the implementation of a condition on $\Theta_{A}(x^{a})$. For a derivation of the hyper-Hermitian metric satisfying a generalization of the first heavenly equation, see [59]. In the present article, we carry out a systematic study of Weyl’s double copy for the Hyper-Hermitian metrics focusing on the family of solutions found in ref. [57]. One of our main results of our work is that for the mentioned family of solutions, the Weyl double copy correspondence implies that one of the two Maxwell’s fields in the single copy has a source current and the other remains sourceless. Such a current is in general, non-conserved.
@@
The paper is organized as follows: in Section 2 we give a brief overview of the double copy for the self-dual case. We describe first the self-dual sector with the kinematic factor taking values in the Lie algebra area preserving diffeomorphisms and its eventual generalization to the kinematical algebra consisting of the Lie algebra of smooth vector fields for the hyper-Hermitian case. In Section 3, after a brief overview on the Weyl double copy, we carry out a systematic study of the Weyl double copy for the hyper-Hermitian metrics. We will focus on the family of metrics called elementary states introduced in [57]. In Section 4 we provide explicit example of solutions of the hyper-Hermitian double copy in the language of the Weyl spinors. Finally in Section 5 is devoted to give our final remarks.
@@ 2 Brief Overview of the Classical Double Copy
In the present section, we briefly overview the double copy procedure between the self-dual sector of the four-dimensional Yang-Mills theory in Euclidean signature and self-dual gravity. Our aim is not to provide an extensive description but to serve to introduce notation and conventions that we will follow in the subsequent sections.
@@ 2.1 Hyper-Kähler double copy
It is known that the double copy has an exact realization in terms of the equation of motion of self-dual Yang-Mills theory starting from a Lagrangian with cubic interactions [36, 37]. In this case, the kinematic algebra is known to be precisely the area-preserving diffeomorphism algebra $\mathfrak{sdiff}$. The Lagrangian formulation used to show this explicitly was based on the description of Leznov and Parkes in which the dynamics of the self-dual sector of Yang-Mills theory is governed by a Lagrangian for a scalar field with a cubic interaction term [40, 41], whose action is given in Euclidean signature by
@@
for a Lie algebra-valued scalar field $\phi(x)=\phi^{a}(x)T_{a}$ on a $4$-dimensional space, where $T_{a}$ are anti-hermitian matrices elements of the Lie algebra $\mathfrak{g}$ and $a=1,\cdots,{\rm dim}\ \mathfrak{g}$. The equations of motion of this action are written as
@@
where $g$ is the Yang-Mills coupling constant, $\partial^{2}\equiv\partial_{x}\partial_{u}+\partial_{y}\partial_{v}$, $\phi$ is a function of the local coordinates on $\mathbb{R}^{4}$. In the momentum-space through the use of the Fourier transform we have
@@
where $f^{abc}$ are the structure constants of $\mathfrak{g}$ (in some representation of the gauge group), which fulfills $[T^{a},T^{b}]=if^{abc}T^{c}$. As part of the BCJ duality, it is remarkable to see that there is a choice of the spacetime coordinates where the kinematic structure constants $X(p_{1},p_{2})$ also fulfill a Jacobi identity just as the gauge structure constants do. Thus, from the structure of equation (4) one can see the Lie algebra structure $\mathfrak{sdiff}\times\mathfrak{g}.$
In the double copy for the classical equations of motion, the second heavenly equation of Plebański is favored with respect to the first one. As we reviewed for the Yang-Mills case (4), there are kinematic factors $X(p_{1},p_{2})$ which satisfies the mentioned kinematic algebra $\mathfrak{sdiff}$ and color factors $f^{abc}$. The double copy correspondence interchanges the color factors with other kinematic factors $f^{abc}\leftrightarrow X(p_{1},p_{2})$ and the gauge and gravity coupling constants $g\leftrightarrow\kappa$. Thus we obtain in momentum space the gravity equation of motion
@@
for a scalar field $\theta$. We now identify the coupling constant as the gravitational coupling $\kappa=\sqrt{32\pi G_{N}}$, in terms of Newton’s constant $G_{N}$. In coordinate space, eq. (5) is simply the Plebański equation for self-dual gravity, also known as the (second) heavenly equation [58]:
@@
where the Poisson bracket in (6) is given with respect the variables $x$ and $y$. Thus, the double copy in the self-dual sector of Yang-Mills theory is particularly transparent and amounts to simply replacing the color algebra with its kinematic counterpart, which is the area-preserving diffeomorphism algebra $\mathfrak{sdiff}$, thus the structure is given by $\mathfrak{sdiff}\times\mathfrak{sdiff}.$
Equation (6) reproduces Plebanski’s second heavenly equation in Euclidean signature [58]:
@@
where we have taken $\kappa=1$. A class of important solutions are obtained when we take equal to zero both terms in (6)
@@
The metric associated with (7) is given by [58]
A solution of (8) is the Sparling-Tod metric [60]
@@
this metric has precisely the form of (9).
@@ 2.2 Hyper-Hermitian double copy
Now we give a brief overview of the hyper-Hermitian spaces (for more details, see [53, 57]). We begin our exploration with a four-dimensional Riemannian manifold denoted as $(\mathcal{M},g)$, where $g$ represents a metric that can be either Lorentzian or Euclidean on $\mathcal{M}$. The metric $g$ is hyper-Hermitian if it is Hermitian with respect to each one of the 3 complex structures $I$, $J$ and $K$ on the manifolds ${\cal M}$, such that they satisfy the relations: $IJ=-JI=K$ plus cyclic permutations of $I$, $J$ and $K$. In this case, the metric $g$ is of the form
@@
where $\Theta^{A}=(\Theta^{0},\Theta^{1})$ is a pair of complex valued functions (or potentials) on ${\cal M}$ which satisfy the Dunajski heavenly equations [57]
@@
where $x^{A}=\{y,-x\}$ and $\partial_{A}\equiv\partial_{x^{A}}$. It is shown in [57] that the metric (11)is a hyper-Hermitian metric and that every hyper-Hermitian metric locally arises under this construction. Within the formalism, it is shown a relation with the hyper-Kähler case described by Plebanski’s second heavenly equation (7). Such relation is given by the existence of a potential $\Theta^{A}$ such that $\nabla_{A0^{\prime}}\Theta^{A}=0$, which in turn, ensures locally the existence of a scalar function $\theta$ which satisfies (7). Every hyper-Hermitian manifold admits a Lee form, which is a one-form $A$ depending on the manifold metric [53, 57]. In this context, Lee form is given in terms of the potential $\Theta^{A}$ in the following form
@@
Thus in momentum space, eqs. (12) implies the integral equations
@@
where $Y^{A}(p_{1C},p_{2B})\equiv p_{2B}\delta^{A}_{C}-p_{1C}\delta^{A}_{B}$ is a kinematic factor which can be proved [44] they satisfy the anti-symmetric property and the Jacobi identity of the Lie algebra of diffeomorphisms of a compact surface $\mathfrak{diff}$. These constitute the Lie algebra of smooth vector fields of a two-dimensional manifold. They are not Hamiltonian vector fields as they do not satisfy the volume preserving condition as was discussed in [57]. As it was elaborated in [44] expression (14) can be regarded as the double copy of eq. (4) via the BCJ mapping $f^{abc}\rightarrow Y^{A}(p_{1C},p_{2B})$. This leads to a $\mathfrak{sdiff}\;\times\;\mathfrak{diff}$ theory.
Equation (12) can be rewritten in coordinates $(x,y,u,v)$ as follows:
@@
A class of solutions to (12) are called hyper-Hermitian elementary states and they satisfy the following equations (which are the hyper-Hermitian analogous of equations (8))
@@
In the subsequent sections, we will study the double copy correspondence associated with these elementary states of the form
@@
where $F_{A}$ is a pair of complex valued functions with explicit dependence according to their arguments.

### v2

@@ Abstract
The self-dual double copy is further explored. In previous work, it has been shown that hyper-Hermitian manifolds also have associated the self-dual gauge theories via Kerr-Schild double copy. The self-dual double copy is generalized in the structure of the kinematic algebra by replacing the area-preserving diffeomorphisms algebra with the diffeomorphisms on a surface algebra. This gave rise to the hyper-Hermitian double copy in the Kerr-Schild approach. In the present article, we further study the hyper-Hermitian case using the Weyl double copy formalism. In particular, we have found solutions within this formalism for different hyper-Hermitian metrics. One of the main features is that there will be two Maxwell spinors and one of them is source-free while the other has a source current. This is compatible with the fact that, in general, the hyper-Hermitian spaces are not Ricci-flat.
@@ 1 Introduction
In the past, the interplay between different theoretical physics subjects has been a way of understanding physical phenomena and is a road to advancing in physics. For instance, the AdS/CFT (or gauge/gravity) correspondence is a duality which (in its more traditional version) connects strong coupling supersymmetric quantum conformal gauge theory in dimension $d$ with classical gravity theory in $d+1$ dimensions (see for instance, ref. [1]). Another relation between gauge and gravitational theories is the one of the double copy and the Bern-Carrasco-Johansson (BCJ) duality [2, 3, 4] which relates scattering amplitudes of non-abelian gauge theories in $d$ dimensions and scattering amplitudes of gravitational theories in the same dimension $d$ (for a complete review, see ref. [5, 6]). The tree-level version has been proved and the loop extension is still at the level of conjecture.
@@
The ideas of the double copy and the BCJ duality have been applied at the level of equations of motion in gauge and gravity theories. Thus it is promoted to the level of solutions of gauge and gravity theories. There is a correspondence between solutions of the equation of motion in both theories [7]. The original version of the classical double copy uses the Kerr-Schild ansatz which implies the existence of certain coordinate systems in which the metric of a gravity theory can be expressed as $g_{\mu\nu}=\eta_{\mu\nu}+\phi k_{\mu}k_{\nu}$. Many applications in diverse field theories were studied for this classical double copy approach, for instance, see refs. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Later another possibility to study the correspondence was through the use of Weyl spinors, which was named the Weyl double copy [25]. The Weyl double copy is reduced to the Kerr-Schild double copy in type D spacetimes [26]. The Weyl double copy relates the curvature of gravity theory written in terms of Weyl spinors to the field strength spinor of a gauge field. After its proposal, Weyl double copy was discussed in different contexts in refs. [27, 28, 29, 30, 31]. This was the precursor to finding a more general description of the double copy in terms of twistor theory [32, 33, 34, 35, 36, 37, 38]. One of the successes of the double copy has been its application as an efficient tool for computing processes involving gravitational waves [29].
@@
The case for the self-dual sector of the Yang-Mills theories in four dimensions and its correspondence with the self-dual sector of Einstein’s theory was proposed and explored for the first time within the context of the double copy[39, 40, 41]. Integrability anomalies arising in the self-dual sectors of gravity and Yang-Mills theory are studied in ref. [42]. This case is of particular relevance since both gauge and gravitational sectors can be described through a cubic Lagrangian giving rise to trivalent Feynman graphs. In the self-dual gauge case, one has the Leznov-Parks Lagrangian [43, 44] and in the gravitational self-dual sector, we have the Plebański Lagrangian [45, 46]. In ref. [47] new cases for the heavenly double copies were studied. There, integrable and non-integrable deformations of the self-dual double copy and generalizations of it as the hyper-Hermitian case were considered. The double copy has been studied as well in the context of celestial amplitudes [48, 49, 50] and higher spin self-dual theories [51]. Some recent works regarding further developments of the self-dual double copy were discussed in refs. [52, 53, 54, 55].
@@
In the present article, we continue the study, which started in ref. [47] and we carry out a deeper study of the hyper-Hermitian double copy using the Weyl double copy. The hyper-Hermitian metrics have been studied previously in refs. [56, 57, 58, 59, 60]. In ref. [60], hyper-Hermitian complex metrics are determined through two holomorphic functions $\Theta_{A}(x^{a})$ (with $A=0,1$) of the coordinates on a complex four manifold. Moreover $\Theta_{A}(x^{a})$ satisfies a generalization of the Plebański’s second heavenly equation [61] that we term it Dunajski’s heavenly equation. Indeed, this generalized equation can be reduced to the second heavenly equation under the implementation of a condition on $\Theta_{A}(x^{a})$. For a derivation of the hyper-Hermitian metric satisfying a generalization of the first heavenly equation, see ref. [62]. In the present article, we carry out a systematic study of the Weyl double copy for the hyper-Hermitian metrics, focusing on the family of solutions found in ref. [60]. One of the main results of our work is that, for the mentioned family of solutions, the Weyl double copy correspondence implies that one of the two Maxwell fields has a source current while the other remains sourceless. Such a current is, in general, non-conserved.
@@
The paper is organized as follows: in Section 2 we give a brief overview of the double copy for the self-dual case. We describe first the self-dual sector with the kinematic factor taking values in the Lie algebra of area-preserving diffeomorphisms and its eventual generalization to the kinematic algebra consisting of the Lie algebra of smooth vector fields for the hyper-Hermitian case. In Section 3, after a brief overview on the Weyl double copy, we carry out a systematic study of the Weyl double copy for the hyper-Hermitian metrics. We will focus on the family of metrics called elementary states introduced in ref. [60]. In Section 4 we provide explicit examples of solutions of the hyper-Hermitian double copy in the language of Weyl spinors. Finally, Section 5 is devoted to give our final remarks.
@@ 2 Brief Overview of the Classical Double Copy
In the present section, we briefly overview the double copy procedure between the self-dual sector of the four-dimensional Yang-Mills theory and self-dual gravity in Euclidean signature. Our aim is not to provide an extensive description but to serve to introduce notation and conventions that we will follow in the subsequent sections.
@@ 2.1 Hyper-Kähler double copy
It is known that the double copy has a realization in terms of the equation of motion of self-dual Yang-Mills theory starting from a Lagrangian with cubic interactions [39, 40]. In this case, the kinematic algebra is known to be precisely the area-preserving diffeomorphism algebra $\mathfrak{sdiff}$. The Lagrangian formulation used to show this explicitly was based on the description of Leznov and Parkes in which the dynamics of the self-dual sector of Yang-Mills theory is governed by a Lagrangian for a scalar field with a cubic interaction term [43, 44], whose action is given in Euclidean signature by
@@
for a Lie algebra-valued scalar field $\phi(x)=\phi^{a}(x)T_{a}$ on a $4$-dimensional space, where $T_{a}$ are anti-hermitian matrix elements of the Lie algebra $\mathfrak{g}$ and $a=1,\cdots,{\rm dim}\ \mathfrak{g}$. The equations of motion of this action are written as
@@
where $g$ is the Yang-Mills coupling constant, $\partial^{2}\equiv\partial_{x}\partial_{u}+\partial_{y}\partial_{v}$ and $\phi$ is a function of the local coordinates on $\mathbb{R}^{4}$. In the momentum-space, through the use of the Fourier transform, we have
@@
where $f^{abc}$ are the structure constants of $\mathfrak{g}$ (in some representation of the gauge group), which fulfils $[T^{a},T^{b}]=if^{abc}T^{c}$. As part of the BCJ duality, it is remarkable to see that there is a choice of the spacetime coordinates where the kinematic structure constants $X(p_{1},p_{2})$ also fulfil a Jacobi identity just as the gauge structure constants do. Thus, from the structure of eq. (4) one can see the Lie algebra structure $\mathfrak{sdiff}\times\mathfrak{g}.$
In the double copy for the classical equations of motion, the second heavenly equation of Plebański is favoured with respect to the first one. As we reviewed for the Yang-Mills case, eq. (4), there are kinematic factors $X(p_{1},p_{2})$ which satisfies the mentioned kinematic algebra $\mathfrak{sdiff}$ and color factors $f^{abc}$. The double copy correspondence interchanges the color factors with another kinematic factor $f^{abc}\leftrightarrow X(p_{1},p_{2})$ and the gauge and gravity coupling constants $g\leftrightarrow\kappa$. Thus, we obtain the gravity equation of motion in momentum space
@@
for a scalar field $\theta$. We now identify the coupling constant as the gravitational coupling $\kappa=\sqrt{32\pi G_{N}}$, where $G_{N}$ is Newton’s constant. In coordinate space, eq. (5) is simply the Plebański equation for self-dual gravity, also known as the (second) heavenly equation [61]:
@@
where the Poisson bracket in eq. (6) is given with respect to the variables $x$ and $y$. Thus, the double copy in the self-dual sector of Yang-Mills theory is particularly transparent, and amounts to simply replacing the color algebra with its kinematic counterpart, which is the area-preserving diffeomorphism algebra $\mathfrak{sdiff}$. Thus, the structure is given by $\mathfrak{sdiff}\times\mathfrak{sdiff}.$
Eq. (6) reproduces Plebanski’s second heavenly equation in Euclidean signature [61]:
@@
where we have taken $\kappa=1$. The metric associated with eq. (7) is given by [61]
A class of important solutions is obtained when we take both terms equal to zero in eq. (6)
@@
A solution of eq. (9) is the Sparling-Tod metric [63]
@@
this metric has precisely the form of eq. (8).
@@ 2.2 Hyper-Hermitian double copy
Now we give a brief overview of hyper-Hermitian spaces (for more details, see refs. [56, 60]). We begin our exploration with a four-dimensional Riemannian manifold denoted as $(\mathcal{M},g)$, where $g$ represents a metric that can be either Lorentzian or Euclidean on $\mathcal{M}$. The metric $g$ is hyper-Hermitian if it is Hermitian with respect to each one of the 3 complex structures $I$, $J$, and $K$ on the manifolds ${\cal M}$, such that they satisfy the relations: $IJ=-JI=K$ plus cyclic permutations of $I$, $J$ and $K$. In this case, the metric $g$ is of the form
@@
where $\Theta^{A}=(\Theta^{0},\Theta^{1})$ is a pair of complex-valued functions (or potentials) on ${\cal M}$ which satisfy the Dunajski’s heavenly equations [60]
@@
where $x^{A}=\{y,-x\}$ and $\partial_{A}\equiv\partial_{x^{A}}$. It is shown in [60] that the metric of eq. (11) is a hyper-Hermitian metric and every hyper-Hermitian metric locally arises under this construction. Within the formalism, a relation with the hyper-Kähler case described by Plebanski’s second heavenly eq. (7) is shown. Such a relation is given by the existence of a potential $\Theta^{A}$ such that $\nabla_{A0^{\prime}}\Theta^{A}=0$, which in turn ensures locally the existence of a scalar function $\theta$ that satisfies eq. (7). Every hyper-Hermitian manifold admits a Lee form, which is a one-form $A$ depending on the manifold metric [56, 60]. In this context, the Lee form is given in terms of the potential $\Theta^{A}$ in the following form
@@
Thus, in momentum space, eq.  (12) implies the integral equations
@@
where $Y^{A}(p_{1C},p_{2B})\equiv p_{2B}\delta^{A}_{C}-p_{1C}\delta^{A}_{B}$ is a kinematic factor. It can be proved [47] that they satisfy the anti-symmetric property and the Jacobi identity of the Lie algebra of diffeomorphisms of a compact surface $\mathfrak{diff}$. These constitute the Lie algebra of smooth vector fields on a two-dimensional manifold. They are not Hamiltonian vector fields, as they do not satisfy the volume-preserving condition as was discussed in ref. [60]. As it was elaborated in ref. [47], eq. (14) can be regarded as the double copy of eq. (4) via the BCJ mapping $f^{abc}\rightarrow Y^{A}(p_{1C},p_{2B})$. This leads to a $\mathfrak{sdiff}\;\times\;\mathfrak{diff}$ theory.
Eq. (12) can be rewritten in coordinates $(x,y,u,v)$ as follows
@@
A class of solutions to eq. (12) is called the hyper-Hermitian elementary states, and they satisfy the following equations (which are the hyper-Hermitian analogous of eq. (9))
@@
In the subsequent sections, we will study the double-copy correspondence associated with these elementary states, which are of
@@
where $F_{A}$ is a pair of complex-valued functions with explicit dependence on their arguments.